OpenStudy (alexandervonhumboldt2):

Top 5 methods of proving Pyth Theorem Give medals if it is helpful!

OpenStudy (alexandervonhumboldt2):

At this point in the scientific literature recorded 367 proof of this theorem. Probably the Pythagorean theorem is the only theorem with such an impressive body of evidence. This diversity can be explained only by the fundamental value theorem in geometry.

OpenStudy (alexandervonhumboldt2):

Through similar triangles: Let ABC has a right-angled triangle with a right angle C. Draw the altitude from C and denote its base by H. triangle ACH is similar to triangle ABC in two corners. Similarly, the triangle CBH is similar to ABC. Introducing the notation: BC=a Ac=b AB=c. ⇒ a/c=HB/a ; b/c=AH/b ⇒ a²=c*HB ; b²=c*AH. ⇒ a²+b²=c*(HB+AH)=c² or c²=a²+b².

OpenStudy (alexandervonhumboldt2):

Proof by equicomplementability There are four equal right-angled triangles, as shown in the figure. Quadrilateral with sides c is a square, as the sum of the two acute angles of 90 °, and the staright angle - 180 °. The total area of the figure is on the one hand, the area of a square with sides (a + b), and on the other hand, the sum of areas of four triangles and squares inner square. (a+b)=4* (ab/2)+c² ⇒ a²+2ab+b²=2ab+c² ⇒ c²=a²+b².

OpenStudy (alexandervonhumboldt2):

Euclid's proof Let A, B, C be the vertices of a right triangle, with a right angle at A. Drop a perpendicular from A to the side opposite the hypotenuse in the square on the hypotenuse. That line divides the square on the hypotenuse into two rectangles, each having the same area as one of the two squares on the legs. Let ACB be a right-angled triangle with right angle CAB. On each of the sides BC, AB, and CA, squares are drawn, CBDE, BAGF, and ACIH, in that order. The construction of squares requires the immediately preceding theorems in Euclid, and depends upon the parallel postulate. From A, draw a line parallel to BD and CE. It will perpendicularly intersect BC and DE at K and L, Connect CF and AD, to form the triangles BCF and BDA. Angles CAB and BAG are both right angles; therefore C, A, and G are collinear. Similarly for B, A, H + angles CBD and FBA are both right angles; therefore angle ABD equals angle FBC, since both are the sum of a right angle and angle ABC. Since AB is equal to FB and BD is equal to BC, triangle ABD must be congruent to triangle FBC. A-K-L is a straight line, parallel to BD, then rectangle BDLK has twice the area of triangle ABD because they share the base BD and have the same altitude BK, Since C is collinear with A and G, square BAGF must be twice in area to triangle FBC. Therefore rectangle BDLK must have the same area as square BAGF = AB². Similarly, it can be shown that rectangle CKLE must have the same area as square ACIH = AC². Now add results: AB² + AC² = BD*BK + KL*KC. Since BD = KL, BD*BK + KL*KC = BD(BK + KC) = BD*BC ⇒ AB² + AC² = BC², since CBDE is a square. This is a quite long proof and not easy one.

OpenStudy (alexandervonhumboldt2):

Proof of Leonardo da Vinci Segment CI divides ABHJ square into two equal parts so as triangles ABC and JHI equal by construction. Using rotation by 90 degrees counter-clockwise around the point A, we see equality and CAJI DABG. Area of painted figures which is equal to the sum of half the area of small squares built on leg and an area of the original triangle. It is equal to half square large square built on the hypotenuse plus area of the original triangle. Thus, half of the amount areas of small squares is equal to half the size of a large square, and hence the sum of areas of the squares constructed on the legs is equal to the area of a square constructed on the hypotenuse.

OpenStudy (alexandervonhumboldt2):

Proof by infinitesimal da/dc=c/a ; c*dc=a*da ; c dc=a da+b db ; c²=a²+b²+constant a=b=c=0 ⇒ constant=0. ⇒ c²=a²+b².

OpenStudy (alexandervonhumboldt2):

@perl

OpenStudy (alexandervonhumboldt2):

@TheSmartOne @bohotness @sleepyhead314

TheSmartOne (thesmartone):

Good Job :D

TheSmartOne (thesmartone):

*Gave the first medal xD)

OpenStudy (alexandervonhumboldt2):

ерч

OpenStudy (alexandervonhumboldt2):

thx

OpenStudy (alexandervonhumboldt2):

sorry for question marks i was wrtting on russian language

OpenStudy (alexandervonhumboldt2):

forgot to smitch languages

TheSmartOne (thesmartone):

Haha, its ok :)